# LLMAID: Identifying AI Capabilities in Android Apps with LLMs

### v1

@@ 3.4. RQ3: How is the usefulness of the AI service summary generated by LLMAID?
Table 5. Different types of identified AI components.
Data Type
Package API
@@
3.5. RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
Regarding this research question, we seek to classify the specific domains and tasks of the identified AI components in the determined AI apps in RQ2. Table 5 shows the number of AI components in different categories. The majority of the AI components (92.82%, 10285/11081) lie within packages and their corresponding APIs. The second largest category of AI components is model, which refers to on-device AI model files, such as .tflite and .caffemodel. Apart from package names, APIs, and on-device AI models, we observe that some apps rely on HTTP requests (e.g., OpenAI APIs) to access cloud-based AI services, which has rarely been discussed in previous studies (xu2019first; li2022ai). The remaining components exhibiting limited shared patterns are collectively classified under the category of Others.
Figure 7. The distribution of tasks (inner ring) and their corresponding AI domains (outer ring) for the AI components.
With these identified AI components, we first reviewed the analysis results to design a classification scheme that captures a wide range of AI domains and tasks. Due to the large number of AI components, we employ ChatGPT to automatically categorize each AI component into a corresponding domain and task. Fig. 7 presents the distributions of classified domains and tasks of the determined AI components. The area of each colored sector represents its corresponding proportion. At the top level, AI components are classified into six common AI domains, including Computer Vision (54.80%), Data Analysis (26.85%)  [^5] Data analysis and processing denote statistical operations within AI frameworks., Natural Language Processing (13.38%), Others (2.75%), Audio and Speech Processing (1.90%), and Augmented Reality (0.32%). The category Others contains AI components that are difficult to classify clearly. The results show that the dominant AI capabilities are still oriented toward Computer Vision and most AI components target the tasks of Object Detection (45.97%), Text Recognition (17.95%), and Face Recognition (16.77%). The second largest AI domain is Data Analysis, which is occupied mostly by the task of Data Processing (86.69%), such as tensor operations and input/output data parsing.
Table 6. Number of Android apps in different domains.
Domains
Data Analysis
Computer Vision
Natural Language
Processing (NLP)
Audio & Speech
Processing (ASP)
Augmented Reality
Others
@@ 4.1.1. LLM Hallucination
Figure 8. The percentage of obfuscated Java file names and their directory names.
4.1.2. Obfuscations in Android apps
Obfuscation techniques, such as string encryption (e.g., the obfuscated API: <<d.f.e.x.b: void <<init>> (int)>>) and Java reflection, are widely adopted in Android apps to safeguard intellectual property against reverse engineering, as extensively evidenced by existing research (dong2018understanding; glanz2020hidden). We have attempted to replicate some deobfuscation artifacts, such as StringHound (glanz2020hidden) and simplify (guillot2010automatic), to recover the obfuscated items in the Android apps in our study. However, these tools either failed to execute successfully in our environment or produced unsatisfactory results. To evaluate how significantly obfuscations affect the final results, we measured the ratio of obfuscation in Android apps in Dataset1. Fig. 8 shows the percentage of obfuscated Java file names and the directory names they reside in. The median value of the obfuscated Java file names and their corresponding folder names account for 11.94% and 20.23%, respectively. While obfuscation may inevitably introduce some limitations, the substantial portion of unobfuscated implementation details still enables us to extract meaningful insights and supports the reliability of our overall analysis and summarization.
4.2. Implications
4.2.1. On-Device and cloud AI integration in Android apps.
The growing integration of on-device models with cloud-based LLM APIs in Android applications introduces new opportunities and challenges that warrant systematic investigation. While such hybrid architectures promise enhanced proactivity, context-awareness, and personalization, they also raise unresolved issues of reliability, privacy, and ethical responsibility. Future research should extend LLMAID to analyze how Android apps orchestrate local and cloud intelligence, assess resilience under failures or connectivity disruptions, and examine the privacy–utility trade-offs inherent in offloading user data to remote services. Equally important is understanding the evolution of developer practices in adopting hybrid AI features, from SDK/API usage to prompt design and deployment strategies, as well as addressing transparency and fairness concerns in user-facing AI behaviors. Advancing these directions will be essential for establishing trustworthy, adaptive, and ethically responsible AI integration in the Android ecosystem.
4.2.2. Security mitigation for software artifacts.
@@ 4.2. Implications
4.2.3. Software artifacts analysis for researchers.
@@
4.2.4. Advanced AI identification approaches & questionnaire with app developers.
@@
In the AI app identification process, we found that some self-claimed AI Android apps, despite describing advanced AI features in their Play Store listings, contained no detectable AI components in the released APKs. A possible reason is the use of Android App Bundles (appbundle:2023), which deliver only essential components at the installation time and provide the necessary AI functions in the on-demand parts, preventing full inspection of the implementation details. To enable comprehensive analysis, it is imperative to adopt a new approach capable of harvesting all parts of an Android app, not only the basic ones but also those loaded on demand. Moreover, not all AI function implementations in Android apps are trivial. There may be some non-trivial implementations that are transparent to us. Therefore, a questionnaire with these popular Android app developers would provide us with a few great insights for further app analysis.
5. Related Work
5.1. AI Mobile Apps Study
@@
Xu et al. (xu2019first) first proposed an APK analysis tool, DL Sniffer to empirically investigate the use of deep learning techniques in Android apps. The DL Sniffer dissects the released APK and identifies the utilization of 16 popular deep learning frameworks via two separate modules. The first one recognizes specific strings (e.g., “TF_AllocateTensor”) in the section rodata of native shared libraries (i.e., file ends with “.so”) via some powerful Linux command (readelf:2023; objdump:2023). The other one identifies AI special APIs in DEX binary files by de-compiling the binary file into smali code via (dex2jar:2023). Experiments on the large-scale dataset containing 16,500 Android apps reveal that top apps are AI techniques early adopters, especially photo beauty apps, and the on-device AI is actually lighter than practitioners expect. Sun et al. (sun2021mind) proposed ModelXRay to determine whether on-device AI frameworks are adopted and to check if the adopted framework is protected or encrypted in a large dataset containing 46,753 Android apps. Their experimental results show that 41% Android apps do not protect the adopted DL models, and even some of the protected models still have a weak status, which would result in the leakage of private information and economic losses. Li et al. (li2022ai) conducted an exploratory study and released a tool, AI Discriminator, to analyze the adoption of AI technologies in Android apps. They first refined the rules used in the identification of AI app and successfully recognized 56,682 AI apps from a total of 7,259,232 mobile apps provided on AndroZoo (allix2016androzoo). With the historical versions of apps considered, they concluded that the AI models utilized in Android apps are not updated regularly and are not protected carefully. What’s more, AI apps always request access to various types of sensitive user data. In addition, they analyzed reviews for the determined AI apps and revealed that users always express positive attitudes towards the adoption of advanced AI technologies.
5.2. LLMs in Software Engineering
@@
Feng et al. (feng2023prompting) tried to reproduce Android bugs with the help of LLMs. They proposed a lightweight bug replay tool, AdbGPT, to automatically rerun Android APKs and trigger bugs according to collected bug reports. The AdbGPT first utilizes prompt engineering , working as real app developers, to extract the necessary steps for bug replay from the given bug reports. And then prompts LLMs to run Android apps under the guidance extracted previously. Their extensive experiments reveal that AdbGPT could obtain a higher accuracy in necessary step extraction compared with the state-of-the-art traditional approaches and finally successfully reproduce 81.3% bugs, which is much better than baselines. Their approach demonstrates the effectiveness of LLMs in assisting traditional software research problems. Tian et al. (tian2023chatgpt) focused on the research tasks of code generation, program repair, and code summarization and conducted an exploratory study to investigate performance in the aforementioned practices. They curated datasets from LeetCode 2016-2020 and LeetCode 2022, and constructed prompts to generate code with problem descriptions, repair buggy code implementations, and summarize code intentions. The experimental results not only reveal that ChatGPT is effective in common software practices but also unveil that limitations exist in LLMs, such as long and detailed prompt descriptions that would hinder the effectiveness of ChatGPT. Ma et al. (ma2023scope) conducted a study to touch on the limitations of the interpretability of ChatGPT in three types of software engineering jobs, including understanding syntax, static behaviors, and dynamic behaviors. Performing ChatGPT on different programming languages, the authors concluded that ChatGPT does have the ability to construct AST but it still produces incorrect or fake facts with respect to dynamic semantic understanding. It is necessary for researchers to propose effective methods to verify the correctness of the generated outputs from ChatGPT. Li et al. (li2023hitchhiker) carried out an LLM-assisted static analysis with respect to use-before-initialization (UBI) bugs for C/C++ programs. They proposed an automatic agent that interacts with both a static analysis tool and an LLM. With the state-of-the-art static analysis tool and meticulous design of prompts, the authors could successfully overcome several challenges, such as modeling bugs specific to the problem, which are hard to manage within separate research directions. They conducted experiments on real potential UBI bugs, achieved higher precision and recall, and even identified 13 previously untouched ones, revealing the effectiveness of LLM in assisting software engineering problems.
6. Conclusion
@@
Data Availability: The source code and datasets are all made publicly available in our artifact package via the following link: https://zenodo.org/records/15522834

### v2

@@ 3.4 RQ3: How is the usefulness of the AI service summary generated by LLMAID?
3.5 RQ4: To what extent can LLMAID uncover prevalent AI capabilities and usage patterns in real-world Android Apps?
Table 5: Different types of identified AI components.
Data
Type
Package
API
@@
Figure 7: Detailed tasks and corresponding AI domains of the AI components.
Regarding this research question, we seek to classify the specific domains and tasks of the identified AI components in the determined AI apps in RQ2. Table 5 shows the number of AI components in different categories. The majority of the AI components (92.82%, 10285/11081) lie within packages and their corresponding APIs. The second largest category of AI components is model, which refers to on-device AI model files, such as .tflite and .caffemodel. Apart from package names, APIs, and on-device AI models, we observe that some apps rely on HTTP requests (e.g., OpenAI APIs) to access cloud-based AI services, which has rarely been discussed in previous studies xu2019first; li2022ai. The remaining components exhibiting limited shared patterns are collectively classified under the category of Others.
Table 6: Number of Android apps in different domains. DADA: Data Analysis; CVCV: Computer Vision; NLPNLP: Natural Language Processing; ASPASP: Audio and Speech Processing; ARAR: Augmented Reality.
Domains DA CV NLP ASP AR Others
@@
With these identified AI components, we first reviewed the analysis results to design a classification scheme that captures a wide range of AI domains and tasks. Due to the large number of AI components, we employ ChatGPT to automatically categorize each AI component into a corresponding domain and task. Fig. 7 presents the distributions of classified domains and tasks of the determined AI components. The area of each colored sector represents its corresponding proportion. At the top level, AI components are classified into six common AI domains, including Computer Vision (54.80%), Data Analysis (26.85%), Natural Language Processing (13.38%), Others (2.75%), Audio and Speech Processing (1.90%), and Augmented Reality (0.32%). The category Others contains AI components that are difficult to classify clearly. The results show that the dominant AI capabilities are still oriented toward Computer Vision and most AI components target the tasks of Object Detection (45.97%), Text Recognition (17.95%), and Face Recognition (16.77%). The second largest AI domain is Data Analysis, which is occupied mostly by the task of Data Processing (86.69%), such as tensor operations and input/output data parsing.
@@ 4.1.1 LLM Hallucination
Figure 8: The percentage of obfuscated Java file names and their directory names.
4.1.2 Obfuscations in Android apps
Obfuscation techniques, such as string encryption (e.g., the obfuscated API: <<d.f.e.x.b: void <<init>> (int)>>) and Java reflection, are widely adopted in Android apps to safeguard intellectual property against reverse engineering, as extensively evidenced by existing research dong2018understanding; glanz2020hidden. We have attempted to replicate some deobfuscation artifacts, such as StringHound glanz2020hidden and simplify guillot2010automatic, to recover the obfuscated items in the Android apps in our study. However, these tools either failed to execute successfully in our environment or produced unsatisfactory results. To evaluate how significantly obfuscations affect the final results, we measured the ratio of obfuscation in Android apps in Dataset1. Fig. 8 shows the percentage of obfuscated Java file names and the directory names they reside in. The median value of the obfuscated Java file names and their corresponding folder names account for 11.94% and 20.23%, respectively. While obfuscation may inevitably introduce some limitations, the substantial portion of unobfuscated implementation details still enables us to extract meaningful insights and supports the reliability of our overall analysis and summarization.
4.2 Implications
4.2.1 Security mitigation for software artifacts.
@@ 4.2 Implications
4.2.2 Software artifacts analysis for researchers.
@@
4.2.3 Advanced AI identification approaches & questionnaire with app developers.
@@
In the AI app identification process, we found that some self-claimed AI Android apps, despite describing advanced AI features in their Play Store listings, contained no detectable AI components in the released APKs. A possible reason is the use of Android App Bundles appbundle:2023, which deliver only essential components at the installation time and provide the necessary AI functions in the on-demand parts, preventing full inspection of the implementation details. To enable comprehensive analysis, it is imperative to adopt a new approach capable of harvesting all parts of an Android app, not only the basic ones but also those loaded on demand. Moreover, not all AI function implementations in Android apps are trivial. There may be some non-trivial implementations that are transparent to us. Therefore, a questionnaire with these popular Android app developers would provide us with a few great insights for further app analysis.
5 Related Work
5.1 AI Mobile Apps Study
@@
Xu et al. xu2019first first proposed an APK analysis tool, DL Sniffer to empirically investigate the use of deep learning techniques in Android apps. The DL Sniffer dissects the released APK and identifies the utilization of 16 popular deep learning frameworks via two separate modules. Experiments on the large-scale dataset containing 16,500 Android apps reveal that top apps are AI techniques early adopters, especially photo beauty apps, and the on-device AI is actually lighter than practitioners expect. In addition, the DL framework developers should take more time to optimize and protect frameworks. Sun et al. sun2021mind proposed ModelXRay to determine whether on-device AI frameworks are adopted and to check if the adopted framework is protected or encrypted in a large dataset containing 46,753 Android apps. Their experimental results show that 41% Android apps do not protect the adopted DL models, and even some of the protected models still have a weak status, which would result in the leakage of private information and economic losses. Li et al. li2022ai conducted an exploratory study and released a tool, AI Discriminator, to analyze the adoption of AI technologies in Android apps. They first refined the rules used in the identification of AI app and successfully recognized 56,682 AI apps from a total of 7,259,232 mobile apps provided on AndroZoo allix2016androzoo. They concluded that the AI models utilized in Android apps are not updated regularly and are not protected carefully. In addition, they analyzed reviews for the determined AI apps and revealed that users always express positive attitudes towards the adoption of advanced AI technologies.
5.2 LLMs in Software Engineering
@@
Feng et al. feng2023prompting tried to reproduce Android bugs with the help of LLMs. They proposed a lightweight bug replay tool, AdbGPT, to automatically rerun Android APKs and trigger bugs according to collected bug reports. Their extensive experiments reveal that AdbGPT could successfully reproduce 81.3% bugs, which is much better than baselines. Their approach demonstrates the effectiveness of LLMs in assisting traditional software research problems. Tian et al. tian2023chatgpt focused on the research tasks of code generation, program repair, and code summarization and conducted an exploratory study to investigate performance in the aforementioned practices. The experimental results not only reveal that ChatGPT is effective in common software practices but also unveil that limitations exist in LLMs, such as long and detailed prompt descriptions that would hinder the effectiveness of ChatGPT. Ma et al. ma2023scope conducted a study to touch on the limitations of the interpretability of ChatGPT in three types of software engineering jobs, including understanding syntax, static behaviors, and dynamic behaviors. Performing ChatGPT on different programming languages, the authors concluded that ChatGPT does have the ability to construct AST but it still produces incorrect or fake facts with respect to dynamic semantic understanding. Li et al. li2023hitchhiker carried out an LLM-assisted static analysis with respect to use-before-initialization (UBI) bugs for C/C++ programs. They conducted experiments on real potential UBI bugs, achieved higher precision and recall, and even identified 13 previously untouched ones, revealing the effectiveness of LLM in assisting software engineering problems.
6 Conclusion
@@
7 Declarations
Funding: Not applicable.
Ethical approval: Not applicable.
Informed consent: Not applicable.
Author Contributions: Pei Liu led the project, oversaw its administration, and drafted the original manuscript. Terry Yue Zhuo developed and implemented the framework. Jiawei Deng curated the dataset and performed the experimental analyses. Zhenchang Xing, Qinghua Lu, Xiaoning Du conceived and designed the study and contributed to manuscript writing. Hongyu Zhang provided critical revisions and editorial improvements.
Data Availability Statements: The source code and datasets are all made publicly available in our artifact package via the following link: https://zenodo.org/records/15958951
Conflict of Interest: The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Clinical trial number: Not applicable.